# Consistency Models for Scalable and Fast Simulation-Based Inference

### v2

@@ 4.2 Experiment 2: Two Moons
This experiment studies the two moons benchmark (Greenberg et al., 2019; Lueckmann et al., 2021; Wiqvist et al., 2021; Radev et al., 2023a). The model is characterized by a bimodal posterior with two separated crescent moons for the observed point $\mathbf{x}_{\text{obs}}=(0,0)^{\top}$ which a posterior approximator needs to recover. We repeat the experiment for different training budgets $M\in\{512,1024,2048,4096,8192\}$ to assess the performance of each method under varying data availability. While $M=512$ is a very small budget for the two moons benchmark, $M=8192$ is considered sufficient.
Results. Trained on a simulation budget of $M=1024$ examples, CMPE consistently explores both crescent moons and successfully captures the local patterns of the posterior (see Figure 2, middle row). Both the affine coupling flow and the neural spline flow fail to fully separate the modes. Most notably, if we aim to achieve fast sampling speed with FMPE by reducing the number of sampling steps during inference, FMPE shows visible overconfidence with 30 sampling steps and markedly deficient approximate posteriors with 10 sampling steps. In stark contrast, CMPE excels in the few-step regime. Figure 4 illustrates that all architectures benefit from a larger training budget. CMPE with 10 steps emerges as the superior architecture in the low- and medium- data regime up to $M=4096$ training instances. Many-step FMPE outperforms the other approximators for the largest training budget of $M=8192$ instances. Keep in mind, however, that many-step FMPE is approximately 30–70$\times$ slower than ACF, NSF, and CMPE in this task.
@@ 4.3 Experiment 3: Inverse Kinematics
Proposed as a benchmarking task for inverse problems by Kruse et al. (2021), the inverse kinematics model aims to reconstruct the configuration $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=\boldsymbol{\theta}\in\mathbb{R}^{4}$ of a multi-jointed 2D arm for a given end position $\mathbf{x}\in\mathbb{R}^{2}$ (see Figure 2, bottom left). The forward process $g:\boldsymbol{\theta}\mapsto\mathbf{x}$ takes as inputs the initial height $\theta_{1}$ of the arm’s base, as well as the angles $\theta_{2},\theta_{3},\theta_{4}$ at its three joints. The inverse problem asks for the posterior distribution $p(\boldsymbol{\theta}\,|\,\mathbf{x})$, which represents all arm configurations $\boldsymbol{\theta}$ that end at the observed 2D position $\mathbf{x}$ of the arm’s end effector. The parameters follow a Gaussian prior $\boldsymbol{\theta}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\sigma}^{2}\mathbf{I})$ with $\boldsymbol{\sigma}^{2}=(\frac{1}{16},\frac{1}{4},\frac{1}{4},\frac{1}{4})$. While the parameter space in this task is 4-dimensional, we can readily plug any parameter estimate $\hat{\boldsymbol{\theta}}$ into the forward simulator $g$ to get the corresponding end effector position $g(\hat{\boldsymbol{\theta}})=\hat{\mathbf{x}}\in\mathbb{R}^{2}$, which is easy to visualize and interpret.
Results. On the challenging small training budget of $M=1024$ training examples, CMPE with $K=30$ sampling steps clearly outperforms all other methods while maintaining fast inference (see Figure 2).
@@ 4.4 Experiment 4: Bayesian Denoising
This experiment demonstrates the feasibility of CMPE for a high-dimensional inverse problem, namely, Bayesian denoising on the Fashion MNIST data set. The unknown parameter $\boldsymbol{\theta}\in\mathbb{R}^{784}$ is the flattened original image, and the observation $\mathbf{x}\in\mathbb{R}^{784}$ is a blurred and flattened version of the crisp image from a simulated noisy camera (Ramesh et al., 2022; Pacchiardi & Dutta, 2022; Radev et al., 2023a).
We compare CMPE and FMPE in a standard and a small data regime. As both methods allow for free-form architectures, we can assess the effect of the architectural choices on the results. Therefore, we evaluate both methods using a suboptimal naïve architecture and the established U-Net (Ronneberger et al., 2015) architecture.
Neural architectures. The naïve architecture consists of a convolutional neural network (CNN; LeCun et al., 2015) to convert the observation into a vector of latent summary statistics. We concatenate input vector, summary statistics, and a time embedding and feed them into a multi-layer perceptron (MLP) with four hidden layers consisting of 2048 units each. This corresponds to a situation where the structure of the observation (i.e., image data) is known, but the structure of the parameters is unknown or does not inform a bespoke network architecture.
In this example, however, we can leverage the prior knowledge that our parameters are images. Specifically, we can incorporate inductive biases into our network architecture by choosing a U-Net architecture which is optimized for image processing (i.e., an adapted version of Nain, 2022). Again, a CNN learns a summary vector of the noisy observation, which is then concatenated with a time embedding into the condition vector for the neural density estimator.
Results. We report the aggregated RMSE, MMD, and the time per sample for both methods and architectures (Table 2). FMPE is not able to generate good samples for the naïve architecture, whereas CMPE produces acceptable samples even in this suboptimal setup. This reduced susceptibility to suboptimal architectures might become a valuable feature in high-dimensional problems where the structure of the parameters cannot be exploited. The U-Net architecture enables good sample quality for both methods, highlighting the advantages of free-form architectures. The MMD values align well with a visual assessment of the sample quality, therefore we deem it an informative metric to compare the results (see Section C.4 for visual inspection). The U-Net architecture paired with a large training set provides detailed and versatile samples of similar quality for CMPE and FMPE. In all cases, CMPE exhibits significantly lower inference time as only two network passes are required for sampling, while achieving better or competitive quality.
@@ 4.5 Experiment 5: Tumor Spheroid Growth
Results. CMPE outperforms the alternative neural methods via better accuracy and calibration, as indexed by lower RMSE and ECE on 400 unseen test instances (see Table 3). The speed of the simpler ACF is unmatched by the other methods. In direct comparison to its free-form contender FMPE, CMPE simultaneously exhibits (i) a slightly higher accuracy; (ii) a drastically improved calibration; and (iii) much faster sampling (see Figure 6 in the Appendix). For this model, FMPE did not achieve satisfactory calibration performance with up to $K=100$ inference steps, so Table 3 reports the best FMPE results for $K=1000$ steps.
@@
5 Conclusion
We presented consistency model posterior estimation (CMPE), a state-of-the-art approach to perform accurate simulation-based Bayesian inference on large-scale models while achieving fast inference speed. CMPE enhances the capabilities of neural posterior estimation by combining the free-form neural architecture property of diffusion models and flow matching with high-performing few-step sampling. To assess the effectiveness of CMPE, we applied it to a set of 3 low-dimensional benchmark tasks that allow intuitive visual inspection, as well as a high-dimensional Bayesian denoising experiment and a scientific tumor growth model. Across all experiments, CMPE consistently outperforms all other methods based on a holistic assessment of posterior accuracy, calibration, and inference speed.
Future work might aim to further reduce the number of sampling steps towards one-step inference, for instance, via extensive automated hyperparameter optimization or tailored training schemes for CMPE. Overall, our results prove the potential of CMPE as a novel simulation-based inference tool, rendering it a new contender for simulation-based inference workflows in science and engineering.
@@ Acknowledgments
We thank Lasse Elsemüller for insightful feedback and thought-provoking ideas. MS acknowledges the funding of the Cyber Valley Research Fund (grant number: CyVy-RF-2021-16) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2075 - 390740016 (the Stuttgart Cluster of Excellence SimTech). MS additionally thanks the Google Cloud Research Credits program and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VP acknowledges funding by the German Academic Exchange Service (DAAD) and the German Federal Ministry of Education and Research (BMBF) through the funding programme “Konrad Zuse Schools of Excellence in Artificial Intelligence” (ELIZA) and support by the state of Baden-Württemberg through bwHPC. VP acknowledges support by the Bundesministerium fuer Wirtschaft und Klimaschutz (BMWK, German Federal Ministry for Economic Affairs and Climate Action) as part of the German government’s 7th energy research program “Innovations for the energy transition” under the 03ETE039I HiBRAIN project (Holistic method of a combined data- and model-based Electrode design supported by artificial intelligence).
Impact statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Consistency training details
Table 4 details all functions and parameters required for training. Except for the choice of parameters $s_{0}$, $s_{1}$, $\varepsilon$, $T_{\text{max}}$ and $\sigma_{\text{data}}$ we exactly follow the design choices proposed by Song & Dhariwal (2023), which improved our results compared to the choices in Song et al. (2023). Song & Dhariwal (2023) showed that that the Exponential Moving Average of the teacher’s parameter $\boldsymbol{\phi}^{-}$ should not be used in Consistency Training, therefore we omitted it in this paper.
Table 4: Design choices: Table and values adapted to our notation from Song & Dhariwal (2023, Table 1).
Metric in consistency loss $d(\mathbf{x},\mathbf{y})=\sqrt{\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}+c^{2}}-c$
Discretization curriculum $N(k)=\min(s_{0}2^{\lfloor k/K^{\prime}\rfloor},s_{1})+1$ where $K^{\prime}=\lfloor K/(\log_{2}\lfloor s_{1}/s_{0}\rfloor+1)\rfloor$
@@
$t_{i}=\left(\varepsilon^{1/\rho}+\frac{i+1}{N(k)-1}\left(T_{\text{max}}^{1/\rho}-\varepsilon^{1/\rho}\right)\right)^{\rho}$, $\rho=7$, $\varepsilon=0.001$, $T_{\text{max}}=200$
@@ B.1 C2ST
Being an expressive classifier, C2ST is most sensitive to regions where there is no overlap between distributions. In our experiments, this results in a pronounced punishment of overconfident posteriors. For limited training data and distributions without sharp edges (e.g., the Gaussian mixture model), some overconfidence is usually given, as samples from the outer tails are simply not present in the training set. As the accuracy is used as a measure, a smaller number of very improbable samples (e.g., samples in the region between two modes) do not influence the results a lot.
In consequence, C2ST may deviate from the visual judgement in important ways. First, overconfident posteriors look more ”precise“, but are punished by C2ST. Second, samples between modes are very salient, but may not influence the results of C2ST a lot.
For distributions with sharp edges and more or less uniform density in a given mode, c2st is very capable in judging the overlap of two distributions, which then is a good measure of quality. This applies to the two-moons example, where we can observe the expected behavior of a larger training budget leading to more accurate results (Figure 4).
@@ B.2 MMD
By capturing and comparing the moments of two distributions, MMD provides a more holistic view than C2ST. It is also applicable to a wider range of cases, as it generalizes to high dimensions. The exact behaviour can be controlled by choosing the kernel function. In our experiments, we observe MMD to be less sensitive for distributions with sharp edges (i.e., two moons), where C2ST seems to provide more reasonable results. We also observe a low sensitivity to samples between modes, leading to unexpected results compared to visual inspection.
@@ B.3 Summary
Calculating meaningful metrics given only samples from two distributions is a challenging task. Ideally, the metrics would reward similar density and punish samples in the low-density regions of the reference distribution. Especially the latter seems to be hard to achieve. Even when metrics work fine for small deviations from the ideal distribution, the behaviour for different forms of larger deviations, as they occur in challenging problems or cases with low training budget, might not be easily grasped or interpreted. That makes it difficult to compare different methods producing qualitatively different deviations (e.g., overconfidence vs. underconfidence) from the reference distribution.

### v3

@@ 4.2 Experiment 2: Two moons
This experiment studies the two moons benchmark [27, 52, 59, 49]. The model is characterized by a bimodal posterior with two separated crescent moons for the observed point $\mathbf{x}_{\text{obs}}=(0,0)^{\top}$ which an approximator needs to recover. We repeat the experiment for different training budgets $M\in\{512,1024,2048,4096,8192\}$ to probe each method under varying data availability. While $M=512$ is a very small budget for the two moons benchmark, $M=8192$ is considered sufficient.
Results. Trained on a simulation budget of $M=1024$ examples, CMPE consistently explores both crescent moons and successfully captures the local patterns of the posterior (see Figure 1, middle row). Both the affine coupling flow and the neural spline flow fail to fully separate the modes. Most notably, if we aim to achieve fast sampling speed with FMPE by reducing the number of sampling steps during inference, FMPE shows visible overconfidence with 30 sampling steps and markedly deficient approximate posteriors with 10 sampling steps. In stark contrast, CMPE excels in the few-step regime. 2(b) illustrates that all architectures benefit from a larger training budget. CMPE with 10 steps emerges as the superior architecture in the low- and medium-data regime up to $M=4096$ training instances. FMPE with 1000 steps outperforms the other approximators by a small margin for the largest training budget of $M=8192$ instances. Keep in mind, however, that 1000-step FMPE is approximately 30–70$\times$ slower than ACF, NSF, and CMPE in this task.
@@ 4.3 Experiment 3: Inverse kinematics
Proposed as a benchmark task for inverse problems by Kruse et al. [53], the inverse kinematics model aims to reconstruct the configuration $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=\boldsymbol{\theta}\in\mathbb{R}^{4}$ of a multi-jointed 2D robot arm for a given end position $\mathbf{x}\in\mathbb{R}^{2}$ (see Figure 1, bottom left). The input to the forward process $g:\boldsymbol{\theta}\mapsto\mathbf{x}$ are the initial height $\theta_{1}$ of the arm’s base, as well as the angles $\theta_{2},\theta_{3},\theta_{4}$ at its three joints. The inverse problem aims to determine the posterior distribution $p(\boldsymbol{\theta}\,|\,\mathbf{x})$, which represents all arm configurations $\boldsymbol{\theta}$ that end at the observed 2D position $\mathbf{x}$ of the arm’s end effector.
Results. On the challenging small training budget of $M=1024$ training examples, CMPE with $K=30$ sampling steps visually outperforms all other methods with respect to posterior predictive performance while maintaining fast inference (see Figure 1). On the C2ST metric, CMPE outperforms normalizing flows (ACF, NSF), but is in turn surpassed by 1000-step FMPE (see 4(b)), which is however 85$\times$ slower (FMPE: 6565ms, CMPE: 78ms). In contrast to the other empirical evaluations, the simulator in this benchmark lacks aleatoric uncertainty [60] and the plausible parameter values are tightly bounded through the geometry of the problem, which might explain this pattern.
Figure 3: Experiment 4. CMPE denoising results on Fashion MNIST (U-Net backbone, $K=2$ sampling steps, 60 000 training images). First row: Original image (target parameters $\boldsymbol{\theta}$). Second row: Blurred image (observations $\mathbf{x}$). Third and fourth row: Means and standard deviations of the approximate posteriors. Note: For standard deviations, darker regions indicate larger variability in the outputs. Adapted from [49]. More in Appendix C.6.
@@ 4.4 Experiment 4: Bayesian denoising
This experiment demonstrates the feasibility of CMPE for a high-dimensional inverse problem, namely, Bayesian denoising on the Fashion MNIST data set [61]. The unknown parameter $\boldsymbol{\theta}\in\mathbb{R}^{784}$ is the flattened original image, and the observation $\mathbf{x}\in\mathbb{R}^{784}$ is a blurred and flattened version of the crisp image from a simulated noisy camera [54, 32, 49]. We compare CMPE and FMPE in a standard (60k training images) and a small data (2k training images) regime. As both methods allow for unconstrained architectures, we can assess the effect of the architectural choices on the results by evaluating both methods using a suboptimal naïve architecture vs. the established U-Net [62].
Neural architectures. The naïve architecture consists of a convolutional neural network [CNN; 63] to convert the observation into a vector of latent summary statistics. We concatenate input vector, summary statistics, and a time embedding and feed them into a multi-layer perceptron (MLP) with four hidden layers consisting of 2048 units each. This corresponds to a situation where the structure of the observation (i.e., image data) is known, but the structure of the parameters is unknown or does not inform a bespoke network architecture. In this example, however, we can leverage the prior knowledge that our parameters are images. Specifically, we can incorporate inductive biases into our network architecture by choosing a U-Net architecture which is optimized for image processing [i.e., an adapted version of 64]. Again, a CNN learns a summary vector of the noisy observation, which is then concatenated with a time embedding into the conditioning vector for the neural density estimator.
@@ 4.5 Experiment 5: Tumor spheroid growth
Results. CMPE outperforms the alternative neural methods via better accuracy and calibration, as indexed by lower RMSE and ECE on 400 unseen test instances (see Table 2). The speed of the simpler ACF is unmatched by the other methods. In direct comparison to its unconstrained contender FMPE, CMPE simultaneously exhibits (i) a slightly higher accuracy; (ii) a drastically improved calibration; and (iii) much faster sampling (see Figure 6 in the Appendix). For this simulator, FMPE did not achieve satisfactory calibration performance with up to $K=100$ inference steps, so Table 2 reports the best FMPE results for $K=1000$ steps.
5 Discussion
We presented consistency model posterior estimation (CMPE), a novel approach to perform accurate simulation-based Bayesian inference on large-scale models while achieving fast inference speed. CMPE enhances the capabilities of state-of-the-art neural posterior estimation by combining few-step sampling with unconstrained neural architectures. To assess the effectiveness of CMPE, we applied it to a set of 3 low-dimensional benchmark tasks that allow intuitive visual inspection, as well as a high-dimensional Bayesian denoising experiment and a scientific tumor growth model. Across our experiments, CMPE emerges as a competitive method for simulation-based Bayesian inference, as evidenced by a holistic assessment of posterior accuracy, calibration, and inference speed.
Limitations. Our proposed CMPE method has two core limitations, which we specifically highlight again in this paragraph. First, consistency models do not directly yield tractable densities for arbitrary parameter values (see Section 3.5 for a remedy). Second, the relation between inference time (number of sampling steps $K$) and performance (posterior accuracy) is not monotonic, as elucidated in Experiment 1, (see 2(a)). We provide more details and report the CMPE posterior quality (quantified by C2ST) as a function of the number of sampling steps for all three benchmark experiments in Section C.5. This phenomenon is not a drawback per se, but rather a counter-intuitive attribute of consistency models. In practice, it can easily be addressed by performing a brief sweep over the number of sampling steps ($K=1,\ldots,K_{\mathrm{max}}$) during inference time and choosing an optimal $K$ based on user-defined metrics (e.g., calibration or MMD on a validation set). Given the very fast inference speed of CMPE and the observed U-shaped relation between sampling steps and performance, this approach does not generally come with a relevant computational burden.
Perspectives. Future work might aim to further reduce the number of sampling steps towards one-step inference in SBI tasks to achieve even faster sampling speed in time-sensitive applications. This might be achieved via extensive automated hyperparameter optimization, tailored training schemes, or consistency model inference in a compressed latent space. Furthermore, consistency trajectory models [CTMs; 66] are a closely related generative model family, whose conditional formulation might prove useful for SBI as well. Overall, our results demonstrate the potential of CMPE as a novel simulation-based inference tool. Its unique combination of highly expressive unconstrained neural networks with fast sampling speed renders CMPE a new contender for SBI workflows in science and engineering, where both time and performance are essential.
@@ Acknowledgments
We thank Lasse Elsemüller for insightful feedback and Maximilian Dax for support with the implementation of the gravitational wave experiment in the Dingo library. MS acknowledges funding from the Cyber Valley Research Fund (grant number: CyVy-RF-2021-16) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2075 - 390740016 (the Stuttgart Cluster of Excellence SimTech). MS additionally thanks the Google Cloud Research Credits program and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VP acknowledges funding by the German Academic Exchange Service (DAAD) and the German Federal Ministry of Education and Research (BMBF) through the funding programme “Konrad Zuse Schools of Excellence in Artificial Intelligence” (ELIZA) and support by the state of Baden-Württemberg through bwHPC. VP acknowledges support by the Bundesministerium fuer Wirtschaft und Klimaschutz (BMWK, German Federal Ministry for Economic Affairs and Climate Action) as part of the German government’s 7th energy research program “Innovations for the energy transition” under the 03ETE039I HiBRAIN project (Holistic method of a combined data- and model-based Electrode design supported by artificial intelligence).
Code
The software code for the experiments is available in a public GitHub repository:
https://github.com/bayesflow-org/consistency-model-posterior-estimation
Additionally, we provide a backend-agnostic implementation of Consistency Model Posterior Estimation (CMPE) in the open source BayesFlow library for amortized Bayesian workflows [55]. Users can choose between a PyTorch, TensorFlow, or JAX backend. BayesFlow is available as open-source software on GitHub (https://github.com/bayesflow-org/bayesflow/) and PyPI.
@@ Appendix A Consistency training details
Table 3 details all functions and parameters required for training. Except for the choice of parameters $s_{0}$, $s_{1}$, $\varepsilon$, $T_{\text{max}}$ and $\sigma_{\text{data}}$ we exactly follow the design choices proposed by Song and Dhariwal [46], which improved our results compared to the choices in Song et al. [18]. Song and Dhariwal [46] showed that that the Exponential Moving Average of the teacher’s parameter $\boldsymbol{\phi}^{-}$ should not be used in Consistency Training, therefore we omitted it in this paper.
Table 3: Design choices: Table and values adapted to our notation from Song and Dhariwal [46, Table 1].
Loss metric $d(\mathbf{x},\mathbf{y})=\sqrt{\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}+c^{2}}-c$
Discretization scheme $N(k)=\min(s_{0}2^{\lfloor k/K^{\prime}\rfloor},s_{1})+1$ where $K^{\prime}=\lfloor K/(\log_{2}\lfloor s_{1}/s_{0}\rfloor+1)\rfloor$
@@
$t_{i}=\left(\varepsilon^{1/\rho}+\frac{i-1}{N(k)-1}\left(T_{\text{max}}^{1/\rho}-\varepsilon^{1/\rho}\right)\right)^{\rho}$, $\rho=7$, $\varepsilon=0.001$, $T_{\text{max}}=200$
@@ B.1 C2ST
As an expressive classifier, C2ST is most sensitive to regions without distribution overlap. In our experiments, this results in a pronounced punishment of overconfident posteriors. For limited training data and distributions without sharp edges (e.g., the Gaussian mixture model), some overconfidence is usually given, as samples from the outer tails are simply not present in the training set. As the accuracy is used as a measure, a smaller number of very improbable samples (e.g., samples in the region between two modes) do not influence the results a lot.
In consequence, C2ST may deviate from the visual judgment in important ways. First, overconfident posteriors look more ”precise“, but are punished by C2ST. Second, samples between modes are very salient, but may not influence the results of C2ST a lot.
For distributions with sharp edges and more or less uniform density in a given mode, c2st is very capable in judging the overlap of two distributions, which then is a good measure of quality. This applies to the two-moons example, where we can observe the expected behavior of a larger training budget leading to more accurate results (2(b)).
@@ B.2 MMD
By capturing and comparing the moments of two distributions, MMD provides a more holistic view than C2ST. It is also applicable to a wider range of cases, as it generalizes to high dimensions. The exact behavior can be controlled by choosing the kernel function. In our experiments, we observe MMD to be less sensitive for distributions with sharp edges (i.e., two moons), where C2ST seems to provide more reasonable results. We also observe a low sensitivity to samples between modes, leading to unexpected results compared to visual inspection.
@@ B.3 Summary
Calculating meaningful metrics given only samples from two distributions is a challenging task. Ideally, the metrics would reward similar density and punish samples in the low-density regions of the reference distribution. Especially the latter seems to be hard to achieve. Even when metrics work fine for small deviations from the ideal distribution, the behavior for different forms of larger deviations, as they occur in challenging problems or cases with low training budget, might not be easily grasped or interpreted. That makes it difficult to compare different methods producing qualitatively different deviations (e.g., overconfidence vs. underconfidence) from the reference distribution.
@@ Appendix C Additional details and results
For neural network training, we used a Mac M1 CPU for Experiments 1–3, an NVIDIA V100 GPU for Experiment 4, an NVIDIA RTX 4090 GPU for Experiment 5, and an NVIDIA H100 GPU for Experiment 6. The evaluation scripts were executed on a Mac M1 CPU for Experiments 1–2, on an NVIDIA V100 GPU for Experiments 3–4, on an NVIDIA RTX 4090 GPU for Experiment 5, and on an NVIDIA V100 GPU for Experiment 6. The computation time for Experiments 1–5 is approximately one day, with the evaluations (e.g., reference posteriors, C2ST computations, MMD computations taking up a considerable amount of time. However, our proposed CMPE method is designed to be carried out on consumer-grade hardware for applied analyses on real data sets, which usually requires minutes to hours of training and can be executed on CPUs depending on the application. In contrast, the large-scale analysis in Experiment 6 requires 2–3 days of training time on high-performance computing infrastructure [see 11, for details on this setting].